YWXZ is a parallelogram with diagonals that intersect at point S. If YS = 2 and SX = x - 7, find x.
Answer by ikleyn(52864)

.
In a parallelogram, diagonals bisect each other.

It implies

2 = x - 7 --> x = 2+7 = 9.